OpenStudy (anonymous):

Miguel is four years older than jasmine. The sum of their ages is twenty eight. How old is jasmine?

OpenStudy (anonymous):

Oh and please write an equation

OpenStudy (anonymous):

m = j + 4 m + j = 28 j + 4 + j = 28 2j = 24 j = 12 Jasmine is 12 years old

OpenStudy (anonymous):

Thanks!